ON MONOCHROMATIC ARITHMETIC

PROGRESSIONS HAVING ODD STEP

VESELIN JUNGIC

Abstract. We perform an experiment concerning certain sets of 2-colorings and the existence of monochromatic arithmetic progressions of odd step. We use the results of the experiment to prove more general statements.